I keep getting two different answers for this problem..Please help.
Find the indicated outputs for:
f(x)= 5x2 -2x
f(0)=
f(-1)=
f(2)=

Answer by stanbon(75887)

Find the indicated outputs for:
f(x)= 5x^2 -2x
f(0)= 5*0^2-2*0 = 0-0 = 0
--------------------------
f(-1)= 5*1^2-2*1 = 5-2 = 3
--------------------------
f(2)= 5*2^2-2*2 = 5*4-4 = 20-4 = 16
=====================================
Cheers,
Stan H.
==========